Research Proposal

Author: Iede Snoek

Title: The Formal Analysis of Opera

Purpose: The construcrtion of a formal theoretical frameworks, in which opera’s can be described, classified, and analyzed

Description: Books about opera always seem to signify the musical side of opera, with few exceptions, like Opera as Drama^{1} by Joseph Kerman.

On the other hand books about theatre seem to emphasize the non-musical theatre, and in those books opera is almost completely neglected. Erika Fischer-Lichte^{2} for the sake of completeness than from a feeling interest.

We can therefore safely state that there is no consistent theoretical framework for opera analysis, as there is for instance in music analysis.

Methodology

A good starting point for the building of such a theoretical framework would be the Aristotelian æsthetics of drama. Aristotle stated that a perfect drama had three unities:

1. Place

2. Time

3. Action

I think those three aspects of drama could be the starting point for our framework. None of these aspects has been thoroughly researched when it comes to opera. In books on non-musical theatre, a kind of theoretical framework can be found, like in Pfister’s book on drama^{3}.

In this thesis I will discuss opera’s from the whole 400 years of opera history. From this epoch I will chose the great composers, to keep things manageable for me. The composers will range from Monteverdi and Peri to Andriessen and Salinen.

1Kerman, Joseph. Opera as Drama, New York 1952 2Fischer-Lichte, Erika Semiotik des Theaters: eine Einführung Teil I. das System der theatralischen Zeichen, Tübbingen, 1983

3Pfister, Manfred The Theory and Analysis of Drama Cambridge 1988